OpenStudy (anonymous):

A pipe is leaking at 1.5 cups per day. About how many gallons per week is the pipe leaking? hint= 1 gallon= 16 cups

OpenStudy (anonymous):

plz help anyone

OpenStudy (anonymous):

ok soooooooo

OpenStudy (anonymous):

multiply your cups given by 16

OpenStudy (tkhunny):

1 gallon = 16 cups 1 week = 7 days That's all you need.

OpenStudy (anonymous):

ok than wat do i multiply them with reacon

OpenStudy (anonymous):

divide that by 7

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

1.5*16=x x/7= your answer

OpenStudy (anonymous):

i got a very big number reacon

OpenStudy (anonymous):

so i have that flipped multiply 1.5*7 and then divide that by 15

OpenStudy (anonymous):

sorry*

OpenStudy (anonymous):

im confussed very

OpenStudy (tkhunny):

Don't be dividing and multiplying willy-nilly. Be consistent and systematic. \(\dfrac{1.5\;cups}{1\;day}\cdot\dfrac{1\;gallon}{16\;cups}\cdot\dfrac{7\;days}{1\;week}\) Find all the units that cancel appropriately and it will tell you which to multiply and which to divide.

OpenStudy (anonymous):

tkyhunny i dont under stand u so u can leave now i need 1 person to help me and only 1

OpenStudy (anonymous):

reacon tell me again wat i have to do

OpenStudy (anonymous):

tkhunny has it right he uses a style called dimensional analysis so you know that you have \[\frac{ 1.5cups }{ 1 day }\] since you are leaking 1.5 cups in each day

OpenStudy (anonymous):

hurry plz can u type faster plz

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

yes sorry lost connection

OpenStudy (anonymous):

oh ok thats fine

OpenStudy (anonymous):

so you want cups on bottom so cups cancel and days on top so days cancel

OpenStudy (anonymous):

so 1.5*7=10.5

OpenStudy (anonymous):

so did u ever find out how to do it

OpenStudy (anonymous):

10.5/16=0.656 gallons/week

OpenStudy (anonymous):

so thats the answer, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thx a lot wat grade do u go to

OpenStudy (anonymous):

um why do you ask?

OpenStudy (anonymous):

idk just wondering

OpenStudy (anonymous):

um id prefer not to answer

OpenStudy (anonymous):

ok it doestnt matter to me no big deal guess wat grade im in

OpenStudy (anonymous):

ok so i ned help with one more question